Find the minimum sample size n needed to estimate μ for the given values of​c,σ​,and E. c=0.90,σ=8.3,and E=2 Assume that a preliminary sample has at least 30 members. n=?
Answer #1

= 1 - 0.90 = 0.10

Z/2 = 1.6449 (From Z table )

Sample size = (Z/2 * / E)2

= ( 1.6449 * 8.3 / 2 )2

= 46.6

Sample size = 47 (Rounded up to nearest integer)